OpenStudy (anonymous):

4/70 ratios simpilest form

OpenStudy (firejay5):

\[\frac{ 2 }{ 35 }\] = \[\frac{ 4 }{ 70 } \div \frac{ 2 }{ 2 } = \frac{ 2 }{ 35 }\]

OpenStudy (firejay5):

Is that what you are looking for?

OpenStudy (anonymous):

help yes

OpenStudy (firejay5):

Need anymore help? @carlos321